Trace the curve y


2 = (x − 1)(x − 2)(x − 3).


1
Expert's answer
2021-01-04T19:42:22-0500

Equation of the curve : "\pmb {y=(x-1)(x-2)(x-3)}"


"y=(x-1)(x-2)(x-3)=x^3-6x^2+11x-6"


Differentiating w.r.t x,

"\frac{dy}{dx}=3x^2-12x+11" (1)


Symmetry :

The curve is not symmetric w.r.t. any line or axis.


Region :

"y=(x-1)(x-2)(x-3)\\"

"y" is defined "\forall x\in \R"

"\therefore" Domain of "y=(-\infty,\infty)"


Now from (1), "\frac{dy}{dx}" exists "\forall x\in \R"

So, "y=f(x)" is continuous "\forall x\in\R" (2)


Now,

"\;\;\;\;\;\;\;\;x\to \infty\Rightarrow y\to \infty\\
and, \;y\to -\infty\Rightarrow y\to-\infty"


So from (2), by intermediate-value theorem,

"\forall y\in \R\;\;\exists x\in\R \ni y=(x-1)(x-2)(x-3)"


"Range \; of \;y=(-\infty,\infty)"

"\therefore x\in(-\infty,\infty),\;\; y\in (-\infty,\infty)"


Point of Intersection :

Putting x=0, we get y=-6

and Putting y=0, we get "x=1,2,3"

"\therefore" The curve cuts x-axis at "(1,0),(2,0),(3,0)" and,

the curve cuts y-axis at "(0,-6)"


Tangent at Origin :

The curve does not go through the origin. So there is no tangent at the origin.


Asymptotes :

Coefficient of highest degree of "x" and "y" are both constant. So, There does not exist any horizontal or vertical asymptote.


Putting "y=mx+c,\\"

"mx+c=x^3-6x^2+11x-6\\
\Rightarrow x^3-6x^2+(11-m)x-(c+6)=0"

"\Rightarrow" Coff. of "x^3=1=Constant"

So There does not exist any oblique asymptote either.


Derivatives :-

"\frac{dy}{dx}=3x^2-12x+11"

Differentiating w.r.t x,

"\frac{d^2y}{dx^2}=6x-12=6(x-2)"


Now

"\frac{dy}{dx}=0\\
\Rightarrow 3x^2-12x+11=0\\
\Rightarrow x=\frac{12\pm \sqrt{12^2-4\times3\times11}}{6}=\frac{6\pm\sqrt3}{3}=2\pm\frac{1}{\sqrt3}"


"\frac{d^2y}{dx^2}_{[x=2+1/\sqrt3]}=\frac{6}{\sqrt3}>0"

"\therefore" There is a local minimum at "x=2+1/\sqrt3"

and,

"\frac{d^2y}{dx^2}_{[x=2-1/\sqrt3]}=-\frac{6}{\sqrt3}<0"

"\therefore" There is a local minimum at "x=2-1/\sqrt3"


Graph :